# Cubic power functions with optimal second-order differential uniformity

### v1

@@ 4. General cubic monomials
Firstly, suppose $\gcd(i,n)\geq 3$. Then there would exist $c\in\mathbb{F}_{2^{\gcd(i,n)}}\subseteq\mathbb{F}_{2^{n}}$, $c\neq 0,1$, such that $c^{2^{i}}+c=0$. Then, from Equation 4.1, we have:
@@
for at most 4 solutions. But there exist at least 8 elements $z\in\mathbb{F}_{2^{\gcd(i,n)}}$, contradicting $f$ being optimal. As such, we must have $\gcd(i,n)\leq 2$. Similarly, by Equation 4.2 we must have $\gcd(j,n)\leq 2$, and by Equation 4.3 we must have $\gcd(j-i,n)\leq 2$. Note then that if $n$ is odd, we have $\gcd(i,n)=\gcd(j,n)=\gcd(j-i,n)=1$. This proves the first statement.
Assume now that $n$ is even, and that $\gcd(i,n)=2$, i.e. $i$ is even also. Again, there would exist $c\in\mathbb{F}_{2^{\gcd(i,n)}}\subseteq\mathbb{F}_{2^{n}}$, $c\neq 0,1$, such that $c^{2^{i}}+c=0$, and from Equations 4.1 and 4.3, we have:
@@ 5. Experimental results
We found exactly two exponents of algebraic degree 4, $d=15,85$ at $n=5,10$ respectively. Observe that these exponents are of the form $d=2^{3k}+2^{2k}+2^{k}+1$ for $k=1,2$. Notably, $d=585$, corresponding to $k=3$ was computed to have second-order differential uniformity 8 for $n=15$.
All other computed optimal exponents are of algebraic degree 3. Every computed optimal cubic exponent was of the form $d=2^{2k}+2^{k}+1$ for some $k$ satisfying $\gcd(k,n)=1$ (up to the aforementioned transformation). In other words, every computed optimal cubic exponent is of the form described in Theorem 3.3.

### v2

@@ 4. General cubic monomials
Firstly, suppose $\gcd(i,n)\geq 3$. Then there would exist $c\in\mathbb{F}_{2^{\gcd(i,n)}}\subseteq\mathbb{F}_{2^{n}}$, $c\neq 0,1$, such that $c^{2^{i}}+c=0$. Then, from Equation (4.1), we have:
@@
for at most 4 solutions. But there exist at least 8 elements $z\in\mathbb{F}_{2^{\gcd(i,n)}}$, contradicting $f$ being optimal. As such, we must have $\gcd(i,n)\leq 2$. Similarly, by Equation (4.2) we must have $\gcd(j,n)\leq 2$, and by Equation (4.3) we must have $\gcd(j-i,n)\leq 2$. Note then that if $n$ is odd, we have $\gcd(i,n)=\gcd(j,n)=\gcd(j-i,n)=1$. This proves the first statement.
Assume now that $n$ is even, and that $\gcd(i,n)=2$, i.e. $i$ is even also. Again, there would exist $c\in\mathbb{F}_{2^{\gcd(i,n)}}\subseteq\mathbb{F}_{2^{n}}$, $c\neq 0,1$, such that $c^{2^{i}}+c=0$, and from Equations (4.1) and (4.3), we have:
@@ 5. Experimental results
All the optimal exponents of algebraic degree 3 were of the form $d=2^{2k}+2^{k}+1$ for some $k$ satisfying $\gcd(k,n)=1$ (up to the aforementioned transformation). In other words, every computed optimal cubic exponent was of the form described in Theorem 3.1. This suggests the following conjecture:
Conjecture 5.1.
[…]
Other than the cubic functions, we found exactly two other optimal exponents, both of them of algebraic degree 4, namely $d=15$ for $n=5$ and $d=85$ for $n=10$. Observe that these exponents are of the form $d=2^{3k}+2^{2k}+2^{k}+1$ for $k=1,2$ respectively. For $k=3$, the exponent of this form would be $d=585$, but this exponent did not exhibit optimal second-order differential uniformity for any $n\leq 20$. It would be of interest to determine a class of optimal power functions of algebraic degree 4 or higher.